A CARNOT ENGINE HAS AN EFFICENCY OF 0.3, WHEN THE TEMPERATURE OF THE SINK IS 350K ,. FIND THE CHANGE IN TEMPERATURE OF THE SOURCE WHEN THE EFFICENCY BECOMES 0.5
 

Grade:12

1 Answers

Arun
6 years ago
Dear student
 
Efficiency (n) = 1 - T2/T1 = 1 - (350/Tsource)
0.3 = 1 - (350/Tsource)
350/ Tsource = 0.7
Tsource = 500 K
Also if (n), efficiency = 0.5
0.5 = 1 - 350/T
350/T = 0.5
T = 700 K
Hence change in temperature of source = 200 K increase
 
Regards
Arun (askIITiajs forum expert)